ganeshie8 (ganeshie8):

If \(p\) is a prime and \(\gcd(k, p-1)=1\), prove that the integers \[\large 1^k,2^k,3^k,\ldots,(p-1)^k \] form a complete set of residues modulo \(p\)

ganeshie8 (ganeshie8):

I have googled but found nothing useful, running low on ideas at the moment... appreciate any kind of favors :)

ganeshie8 (ganeshie8):

basically we need to show that the integers \[\large 1^k,2^k,3^k,\ldots,(p-1)^k\] are congruent in some order to the integers \[1,2,3,\ldots, (p-1)\] in modulo \(p\)

OpenStudy (kainui):

Well I don't know what a residue is, it's like some kind of complex analysis thing that's also a number theory thing... But I might be able to help anyways, here goes. If p wasn't prime, p-1 would still not share any divisors with p so it seems important or related. Specifically gcd(p, p-1)=1 is what I'm saying.

OpenStudy (kainui):

Ok so are you saying that gcd(5,7-1)=1 so then this means we can write 1^6, 2^6,... 5^6 and that set will be all the natural numbers less than 6?

ganeshie8 (ganeshie8):

Yes two consecutive integers are always relatively prime we're given that the exponent \(k\) is relatively prime to \(p-1\)

ganeshie8 (ganeshie8):

small correction : `Ok so are you saying that gcd(5,7-1)=1 so then this means we can write 1^5, 2^5,... 6^5 and that set will be all the natural numbers less than 7?`

OpenStudy (kainui):

Ahh I see, less than the prime itself ok.

ganeshie8 (ganeshie8):

Exactly! 1^5 = 1 mod 7 2^5 = 4 mod 7 3^5 = 5 mod 7 4^5 = 2 mod 7 5^5 = 3 mod 7 6^5 = 6 mod 7 we get the complete set {1, 2, 3, 4, 5, 6} as residues as desired

OpenStudy (usukidoll):

yo wassup does anyone know matlab? kind of stuck :(...ooh number theory

ganeshie8 (ganeshie8):

recently i remember seeing @dan815 @UnkleRhaukus @Zarkon using matlab..

OpenStudy (usukidoll):

damnnn... I need to modify a code but I'm stuck

OpenStudy (kainui):

I'm just trying to break it to see what happens when p-1 has a factor in common. I'm still playing around with it, kind of a weird problem to get my head around.

ganeshie8 (ganeshie8):

Interesting... specifically when \(k\) is a multiple of \(p-1\) we are gettomg all \(1\)'s as residues : ``` 1^6 = 1 mod 7 2^6 = 1 mod 7 3^6 = 1 mod 7 4^6 = 1 mod 7 5^6 = 1 mod 7 6^6 = 1 mod 7 ```

OpenStudy (kainui):

Another interesting pattern, gcd(4, 7-1) = 2 ``` 1^4 = 1 2^4 = 2 3^4 = 4 4^4 = 4 5^4 = 2 6^4 = 1 ``` mod 7 of course. =P

ganeshie8 (ganeshie8):

That looks neat xD may be we can conjecture : \(x^k\equiv a \pmod{p} \) has exactly \(d\) solutions where \(d = \gcd(k, p-1)\)

ganeshie8 (ganeshie8):

need to modify a bit : If \(x^k\equiv a \pmod{p} \) has a solution, then it has exactly \(d\) solutions, where \(d = \gcd(k, p-1)\)

ganeshie8 (ganeshie8):

for exampile : in your table \(x^4\equiv 2 \pmod{7}\) has \(2\) solutions : \(x=2, 5\) because \(\gcd(4, 7-1)=2\)

ganeshie8 (ganeshie8):

So looks like we can generalize the actual problem statement

ganeshie8 (ganeshie8):

i feel that can be complicated to prove compared to our original problem..

OpenStudy (kainui):

I found this sort of thing that's true I think for greater than 2 or 3, might be helpful or related to think about: \[\Large a^n \equiv a \mod (a(a-1))\] Maybe the more general is easier, I don't know I really don't know much about how to think with these things haha.

OpenStudy (kainui):

I'm looking through my abstract algebra book for ideas lol

OpenStudy (kainui):

Why do we need a prime in the first place?

OpenStudy (kainui):

I guess it makes it so that the only possible prime p-1 can be is 2, and maybe that is a special case or something so it works by chance.

ganeshie8 (ganeshie8):

I'm not so sure.. maybe the author wants us to use Fermat's little thm.. but i don't see how it helps here..

ganeshie8 (ganeshie8):

earlier you're showing \(a^n-a\) is divisible by \(a(a-1)\) ? this looks neat, im still trying the proof xD

ganeshie8 (ganeshie8):

\(a^n-a = a(a^{n-1}-1) = a(a-1)(stuff)\) nice :)

OpenStudy (kainui):

Yeah, I feel like I can't stop from seeing the geometric series every single day, and that was it lol.

OpenStudy (kainui):

\[\large \frac{a}{a} \frac{a^{n-1}-1}{a-1} = a^n+a^{n-1}+ \cdots + a+1 \equiv 1 \mod a(a-1)\]

ganeshie8 (ganeshie8):

i like this proof more xD

OpenStudy (kainui):

Honestly I had no idea there was this connection, I just figured this random thing out and never proved it but knew it had to be true cause of how I discovered it. But then what you said made me see it immediately... But it's almost so good looking that I'm not even sure if it's true lol.

OpenStudy (kainui):

That last step seems wrong after I put it, I just sort of assumed that all the powers of a all divided out leaving just 1. But that seems wrong since it should be 0 I think not 1.

OpenStudy (kainui):

At any rate, the proof by induction is simple, start here and then you can see how you could do this n times, kind of like multiplying an idempotent matrix by itself or I guess more closely related to boolean algebra. \[\large a^2 -a \equiv 0 \mod a(a-1)\] Hmm.

ganeshie8 (ganeshie8):

your previous proof works with a slight modification : \[\large \frac{a^{n}-1}{a-1} = a^{n-1}+ \cdots + a+1 \equiv 1 \pmod {a}\] multiply \(a-1\) both sides \[\large a^{n}-1 \equiv a-1 \pmod {a(a-1)}\] \[\large a^{n} \equiv a \pmod {a(a-1)}\] :D

OpenStudy (kainui):

I think I just repeated your argument in trying to find a better way lol. \[\large a(a^{n-1}-1) \frac{a(a-1)}{a(a-1)} \equiv a(a-1) (stuff)\]

ganeshie8 (ganeshie8):

lets pretend we dont know how to factor x^n-1 :P

OpenStudy (kainui):

Yeah we need to use fermats little theorem here I am feeling it

ganeshie8 (ganeshie8):

then we can prove the same using divisibility arguments alone like this : \[a^n-a\equiv 0 \pmod {a}\tag{1}\] \[a^n-a\equiv (a-1+1)^n-(a-1+1) \equiv (0+1)^n - (0+1) \equiv 0 \pmod {a-1}\tag{2}\] the result follows from above two congruences

ganeshie8 (ganeshie8):

since a| (a^n-a) and (a-1) | (a^n-a) we have a(a-1) | (a^n-a)

OpenStudy (kainui):

Oh so in the binomial expansion of (a-1) with 1 to the power n, all the terms go away except 1? Is that how you're thinking of it or is what you're saying kind of different somehow, I just have trouble with these mod things.

ganeshie8 (ganeshie8):

Yes... we're just expanding ( `a-1` + 1)^n using binomial thm..

ganeshie8 (ganeshie8):

Also, we can simply replace `a-1` by `0` because `a-1 = 0 (mod a-1)`

ganeshie8 (ganeshie8):

( `a-1` + 1)^n = ( `0` +1)^n (mod `a-1`)

OpenStudy (kainui):

Oh that's pretty simple I like that better.

OpenStudy (kainui):

Are you familiar with Euler's Theorem? Apparently that's what it's called lol...

OpenStudy (kainui):

a and m are coprime, then \[\large a^{\phi(m)} \equiv 1 \mod m\] It's a generalization of Fermat's theorem basically where \[\large \phi(m)\] is the totient function, which apparently gives the number of coprime numbers less than the number. So for primes \[\large \phi(p)=p-1\] which is how we recover Fermat's little theorem out of it. I think we can play with this to prove your thing.

OpenStudy (usukidoll):

EUler's Theorem :D

ganeshie8 (ganeshie8):

Yes when m is a prime, euler = fermat

OpenStudy (usukidoll):

gawd I love calculating phi, d, sigma, and meu that was the shiz.

ganeshie8 (ganeshie8):

lol i love sigma and tau functions xD

OpenStudy (kainui):

oh quick question before I go to sleep is there anything like this: \[\Large \lim_{n \to \infty} \sum_{n=1}^\infty a_n x^n \mod 2n\] Basically a way to cut out all even numbers from a generating series or something like this as an application.

OpenStudy (kainui):

Wait that doesn't make sense, I'm going to bed, also what's sigma tau and mu functions?

ganeshie8 (ganeshie8):

I'll try this again in my evening... thanks Kai :) \(\large \tau(n)\) : number of positive divisors of \(n\) \(\large \sigma(n)\) : sum of those divisors

OpenStudy (anonymous):

If it helps think of it asinfinite zeta also I remember I have proven this with nt concepts will try and send snap

OpenStudy (anonymous):

Okk I remember it from abstract algebra as well

OpenStudy (anonymous):

Homomorphism to something

OpenStudy (thomas5267):

Anything to do with binomial theorem?

OpenStudy (kainui):

What about Wilson's theorem could we use that?

OpenStudy (kainui):

\[\large (p-1)! \equiv -1 \mod p\] Seems related

OpenStudy (anonymous):

p is prime , let p have a primitive root of \(r\le p-1~and ~gcd(r,p)=1\) then \(1^k,...(p-1)^k\) its congruent to \(r,r^2,r^3,.....,r^{p-1}\) in some order

OpenStudy (anonymous):

are congruent in modulo p * QED

OpenStudy (anonymous):

or if i wanna use ind , let k be the smallest that satisfies \(a\equiv r^k \mod p ,~gcd(r,p)=1\) in other way let k be the index of a relative to r \(ind ~a^k=k ~ind ~a \) hmm directly , right ?

ganeshie8 (ganeshie8):

Ahh that looks neat xD i completely forgot about primitive roots here

OpenStudy (anonymous):

hehe im not perfectly convincing myself

ganeshie8 (ganeshie8):

Here I tried to complete your proof : Let \(r\) be a primitive root of \(p\), then the given integers are congruent in some order to below integers \[\large r^{1k},r^{2k},\ldots,r^{(p-1)k}\] We conclude the proof by showing that none of these powers are congruent to each other. If any of these powers are congruent to each other, for \(i\ne j\): \[r^{ik}\equiv r^{jk}\pmod{p} \] dividing \(r^{jk}\) both sides \[r^{ik-jk}\equiv 1\pmod{p} \] that yields \(ik-jk \equiv 0 \pmod{p-1}\) dividing \(k\) both sides gives \(i-j \equiv 0 \pmod{p-1}\) which is a contradiction. That means the given set of integers(\(1^k,2^k, \ldots,(p-1)^k\)) are incongrunt to each other. Since there are exactly \(p-1\) incongruent integers, these must be congruent to \(1,2,3,\ldots,(p-1)\) in some order. \(\blacksquare\)

OpenStudy (anonymous):

yes exactly i missed out some steps :P